OpenStudy (anonymous):

At Sam's Hot Dog Stand, 350 hot dogs were sold in one day. The next day, the number of hot dogs sold was 385. What was the percentage of increase in the number of hot dogs sold?

OpenStudy (anonymous):

A 10% B 9% C 35% D 110%

OpenStudy (anonymous):

is it d???

OpenStudy (phi):

no. If you sold the same number on each day, the increase would be 0%. You only sold a few more.

OpenStudy (anonymous):

it is c..

OpenStudy (phi):

You do it the same way you did your previous problem. Change/original_value

OpenStudy (anonymous):

350+350x=385

OpenStudy (anonymous):

what do u think the answer is from the options???

OpenStudy (anonymous):

solve for x!

OpenStudy (anonymous):

????????????????

OpenStudy (anonymous):

350+350x=385

OpenStudy (anonymous):

350x=35 x=??

OpenStudy (anonymous):

10??

OpenStudy (anonymous):

x=.1, in percentage, 10%

OpenStudy (phi):

350 hot dogs 385 hot dogs next day change = 385-350 % increase =100* (385-350)/350